OpenStudy (mathstudent55):

The length of a rectangle is 5 units and its width is 4 units. What is the approximate length of the diagonal of the rectangle? 5 units 6.4 units 8.5 units 9 units

OpenStudy (mathstudent55):

@openstudy_girl2

OpenStudy (mathstudent55):

Use the Pythagorean theorem \(a^2 + b^2 = c^2\) \(5 ^2 + 4^2 = c^2\) \(25 + 16 = c^2\) \(c^2 = 41\) \(c = sqrt{41} \approx 6.4\) Answer: 6.4 units